OpenStudy (arbyj):

I NEED HELP! best/right answer will get a medal! (problem will be in the comments!)

OpenStudy (arbyj):



OpenStudy (anonymous):

64.983

OpenStudy (anonymous):

15/7=tan(m) m=tan-1(15/7) m=64.983 aaprromixatly

OpenStudy (arbyj):

@ttb123456789 how bout angle n?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tan n=7/15 n=25.01689348

OpenStudy (anonymous):

there is one more way find the hpotinuse using pyghogorian theorem. then count the cos(n)=15/hypotinuse

OpenStudy (arbyj):

thank you ! (:

OpenStudy (anonymous):

medal please?

OpenStudy (anonymous):

tahnk you